calculate the temperature of a town situated on the top of a mountain at a height of 600m above sea level , given that the sea level temperature is 37 degrees.
1
Expert's answer
2020-10-15T06:36:54-0400

Every 100 m, the temperature decreases for 0.6°C, which means that 600 m mean 6*0.6=3.6°C temperature drop. This gives


"t_\text{up}=t_\text{down}-3.6=37-3.6=33.4°\text{C}"